You want to obtain a sample to estimate a population mean. Based on previous evidence, you believe the population standard deviation is approximately σ = 45.7 . You would like to be 95% confident that your estimate is within 0.1 of the true population mean. How large of a sample size is required? n=
Answer #1

Answer:

C=  95%, E= 0.1 ,  σ = 45.7

formula for sample size is

n= Za/2*0

here {Z_{\alpha / 2} is the Z critical value for C= 95%

We can get critical value using normal z table

{Z_{\alpha / 2} = 1.96

n = 1.96 * 45.7 (- 0.1

n= 802314.3184

n= 802315

Required sample size is = 802315